OpenStudy (anonymous):

18c(cubed)+3c(squared)d-16cd(squared)+4d(cubed)/3c-2d

OpenStudy (king):

Can you please write as a equation....

OpenStudy (saifoo.khan):

\[\frac{18c^3+3c^2d-16cd+4d^3}{3c-2d}\]

OpenStudy (saifoo.khan):

Like that?

OpenStudy (king):

Thnx!!

OpenStudy (anonymous):

yeah but after -16cd, the d is squared

OpenStudy (saifoo.khan):

u wrote cubed. :O

OpenStudy (saifoo.khan):

\[\frac{18c^3+3c^2d-16cd+4d^2}{3c-2d}\]

OpenStudy (anonymous):

-16cd(squared)

OpenStudy (saifoo.khan):

\[\frac{18c^3+3c^2d-16cd^2+4d^3}{3c-2d}\]

OpenStudy (anonymous):

YES!

OpenStudy (saifoo.khan):

\[6 c^2+5 c d-2 d^2 \]

OpenStudy (anonymous):

how did you get that?

OpenStudy (saifoo.khan):

calculator.

OpenStudy (saifoo.khan):

buy factoring will work here.

OpenStudy (anonymous):

so what do u factor?

OpenStudy (anonymous):